A bag contains 3 white, 1 black and 3 red ball. Two balls are drawn from the well shaked bag. Find the probability of both the balls being black and its mean.    



1
Expert's answer
2021-06-01T16:19:04-0400

Because we replace the first ball before selecting the second, the events are independent, so the probabilities are:

P(both the balls being black) "= \frac{1}{7} \times \frac{1}{7} = \frac{1}{49}"

The mean probability is the same "\frac{1}{49}"